# Lyrics: Boosting Fine-grained Language-Vision Alignment via Semantic-aware Visual Objects

### 08 Dec 2023 v1

@@ Abstract
Large Vision Language Models (LVLMs) have demonstrated impressive zero-shot capabilities in various vision-language dialogue scenarios. However, the absence of fine-grained visual object detection hinders the model from understanding the details of images, leading to irreparable visual hallucinations and factual errors. In this paper, we propose Lyrics, a novel multi-modal pre-training and instruction fine-tuning paradigm that bootstraps vision-language alignment from fine-grained cross-modal collaboration. Building on the foundation of BLIP-2, Lyrics infuses local visual features extracted from a visual refiner that includes image tagging, object detection and semantic segmentation modules into the Querying Transformer, while on the text side, the language inputs equip the boundary boxes and tags derived from the visual refiner. We further introduce a two-stage training scheme, in which the pre-training stage bridges the modality gap through explicit and comprehensive vision-language alignment targets. During the instruction fine-tuning stage, we introduce semantic-aware visual feature extraction, a crucial method that enables the model to extract informative features from concrete visual objects. Our approach achieves strong performance on 13 held-out datasets across various vision-language tasks, and demonstrates promising multi-modal understanding and detailed depiction capabilities in real dialogue scenarios.
@@ 1 Introduction
However, widely-used LVLMs habitually adopt Vision Transformer (ViT) [14] from pre-trained CLIP [42] as the image encoder, whose visual feature generalization capabilities are manifested in executing pre-defined label classification and brief image-text matching. Therefore, learning to effectively detect fine-grained visual objects within images (e.g. color, count, detailed description) and capturing visual morphology (e.g. action recognition, localization) present considerable challenges due to the lack of precise local visual features. Therefore, in situations where the image encoder fails to provide sufficient visual signals to meet the requirements of the specific objectives mentioned in the instructions, the LVLMs tends to produce incorrect responses that deviate from the details of the image. More samples are presented and analyzed in Section 5.5.
@@
we conduct extensive experiments on diverse vision-language tasks, including image captioning, visual question answering and referring expression comprehension. The results demonstrate that Lyrics can achieve state-of-the-art or comparable performance on several benchmarks compared to previous LVLMs.
@@ 2.2 Large Vision-Language Models
With remarkable generalization and robustness of LLMs, common Large Vision-Language Models (LVLMs) use a vision-language cross-modal adapter to align the visual features outputted by the visual encoder with the LLMs, thereby stimulating the ability of LLMs to perceive and understand visual signals. Flamingo [2] freezes the pre-trained visual encoder and LLMs and integrates multi-modal representations through perceiver and gated cross-attention, demonstrating impressive few-shot capabilities. Meanwhile, BLIP-2 [25] trains a Q-Former to compress visual features as input to the frozen LLMs. On this basis, InstructBLIP [11] proposes instruction-aware visual feature extraction that enables flexible and informative feature extraction according to the given instructions. Early work such as LLaVA [27] and Mini-GPT4 [63] attempt to simply feed visual features into LLMs using only a learnable fully-connected layer, which introduce visual instruction tuning to enhance instruction following capabilities in LVLMs. Furthermore, recent works such as Vision-LLM [55], Kosmos-2 [39], Shikra [6] and concurrent Qwen-VL [3] also demonstrate that after the open training on visual encoders and LLMs, the LVLMs can understand located objects within the images and generate text formats of bounding boxes to perform visual grounding.
@@ 3.2 Bootstraping Vision-Language Representation Alignment via Multi-Task Pre-training
(ITM) is a binary classification task that transmits the pooled query embedding and into a classifier followed by softmax to predict a two-class probability and . It aims to learn multi-modal image-text representations to capture both coarse-grained and fine-grained semantic alignments between vision and language. We employ a regionalized bi-directional self-attention mask that permits mutual interaction between visual queries and image caption, as well as between grounding queries and spatial representation, while remaining tokens are prevented from attaching to each other. Let and denote 2-dimension one-hot vectors representing the ground-truth label. The ITM loss is:
@@
The full pre-training objective of MQ-Former can be formulated as:
Model | Image Caption | General VQA | ||||||||||||||
VQAv2 | OKVQA | GQA |
Flamingo-9B | 79.4 | - | 61.5 | 51.8 | 44.7 | - | - | 28.8 | ||||||||
Flamingo-80B | 84.3 | - | 67.2 | 56.3 | 50.6 | - | - | 31.6 | ||||||||
IDEFICS-9B (LLaMA-7B) | 46.0 | 36.8 | 27.3 | 50.9 | 38.4 | - | 44.2 | 35.5 | ||||||||
IDEFICS-80B (LLaMA-65B) | 91.8 | 65.0 | 53.7 | 60.0 | 45.2 | - | 68.9 | 36.0 | ||||||||
BLIP-2 (Vicuna-13B) | - | 103.9 | 71.6 | 65.0 | 45.9 | 41.0 | 61.0 | 19.6 | ||||||||
InstructBLIP (Vicuna-13B) | - | 121.9 | 82.8 | - | - | 49.5 | 63.1 | 33.4 | ||||||||
Shikra (Vicuna-13B) | 117.5 | - | 73.9 | 77.4 | 47.2 | - | - | - | ||||||||
Qwen-VL (Qwen-7B) | - | 120.2 | 81.0 | 78.2 | 56.6 | 57.5 | 68.2 | 38.9 | ||||||||
Lyrics (Vicuna-13B) | 121.1 | 126.8 | 85.4 | 81.2 | 58.2 | 62.4 | 71.1 | 37.6 |
Model | TextVQA | OCR-VQA |
BLIP-2 (Vicuna-13B) | 42.4 | - |
InstructBLIP (Vicuna-13B) | 50.7 | - |
Qwen-VL (Qwen-7B) | 61.5 | 70.5 |
PALI-X-55B | 71.4 | 75.0 |
Lyrics (Vicuna-13B) | 69.4 | 75.8 |
@@ 4 Experiment Setting
Training Data
In the pre-training stage, we use a large-scale, web-crawled set of image-text pairs and filter out low-relevant samples, including LAION [45], CC12M [5], CC3M [47] and SBU [38]. In the instruction fine-tuning stage, we introduce high-quality vision-language annotation and instruction-response data [26, 27, 62, 57] to enhance the instruction following and dialogue capabilities of Lyrics. we use a wide range of publicly available vision-language datasets and transform them into instruction fine-tuning format, covering 6 tasks and 24 datasets. We present detailed description and statistics for each dataset, instructions for training and generating in Appendix.
Implementation Detail
For model settings, we choose ViT-L/14 [14] initialized from pre-trained CLIP [42] via contrastive learning as the image encoder. We build the visual refiner by combining Grounding-DINO-T [59] with 900 output object boxes and Swin-T backbone, SAM-HQ [22] with MAE and pre-trained VIT-H image encoder, and RAM++ [60] with Swin-B backbone. We use Vicuna-13B [9], an instruction-tuned variant from LLaMA [51], as the foundation backbone. Throughout the entire training process, the image encoder and visual refiner remains frozen. We focus on training the MQ-Former and linear projection layer, and efficient fine-tuning the large language model using LoRA [19]. With LoRA, we fine-tune the and via low-rank adaptation. We use images of size 224×224, augmented with random resized cropping and horizontal flipping. Our model has a total of 14.9 billion parameters, of which 278 million are trainable parameters, including the MQ-Former (195M), the projection layer (4M) and LLM with LoRA training (79M).
We use AdamW [28] optimizer with , and a weight decay of 0.05. We use a cosine learning rate scheduler to train our model decay with a peak learning rate of 1e-4 and a linear warmup ratio of 15%. We train the Lyrics on 16xA100 GPUs for 600,000 steps in the vision-language representation alignment stage with a global batch size of 512, and 200,000 steps in the vision-to-language generative learning stage with a global batch size of 64.
@@
Model | GQA | SciQA-Img | VizWiz | TextVQA | RefCOCOg | |
val | test | |||||
Lyrics (Vicuna-13B) | 62.4 | 71.1 | 37.6 | 69.4 | 87.23 | 88.26 |
w/o Grounding Quries | 58.1 (4.3) | 64.3 (6.8) | 34.4 (3.2) | 61.1 (8.3) | 83.10 (4.13) | 83.04 (5.22) |
w/o Spatial Representation | 57.3 (5.3) | 63.7 (7.4) | 34.9 (2.7) | 63.2 (6.2) | 82.66 (4.57) | 83.95 (4.31) |
w/o pre-training stage | 38.3 (24.1) | 28.2 (42.9) | 14.7 (22.7) | 29.3 (40.1) | 31.55 (55.68) | 36.26 (52.00) |
@@ 5 Experiment Result
In this section, we conduct a comprehensive evaluation across various multi-modal tasks to thoroughly assess the visual understanding and generating capabilities of Lyrics, and compare our methods with the state-of-the-art visual-centric generalist models under zero-shot and few-shot settings, primarily including Flamingo [2], IDEFICS [24], BLIP-2 [25], InstructBLIP [11], Shikra [6], Qwen-VL [3] and task-specific methods.
@@ 5.1 Dataset and Evaluation Metrics
We evaluate our model across a range of image captioning, visual question answering (VQA) and referring expression comprehension (REC) benchmarks. For image captioning, we choose COCO [7], Nocaps [1] and Flickr30K [40] as benchmarks and report CIDEr score [53] as metric. We consider five benchmarks including VQAv2 [17], OKVQA [31], GQA [20], ScienceQA (Image Set) [29] and Vizwiz [18] benchmarks for general VQA, two benchmarks including TextVQA [48] and OCR-VQA [34] for text-oriented VQA, and evaluate the performance by matching the model’s response to the ground-truth and reporting top-1 accuracy. We use a sort of REC benchmarks such as RefCOCO [21], RefCOCO+ [30] and RefCOCOg [30] to verify the image understanding and localization capabilities of our models. A predicted bounding box is considered as correct for reporting accuracy if its IOU between prediction and ground-truth is higher than 0.5. We use an open-ended approach with a greedy decoding strategy.
5.2 Image Captioning and Visual Question Answering Results
We first evaluate Lyrics on multiple image captioning and general VQA benchmarks. As demonstrated in Table 1, We discover that Lyrics achieves the best performance across 7 out of 8 benchmarks, and demonstrate competitive results in the remaining VizWiz. Lyrics consistently surpasses its original backbone BLIP-2 by a significant margin across all benchmarks, and achieves competitive performance to Qwen-VL, which possesses a more robust LLM backbone and underwent more pre-training and instruction fine-tuning steps. For instance, we achieve the 121.1, 126.8 and 85.4 state-of-the-art CIDEr scores on three image captioning benchmarks, even outperforms previous generalist models with much more parameters (e.g., Flamingo and IDEFICS with 80B parameters). We achieve 62.1 average accuracy on all benchmarks for general VQA tasks, representing a relative improvement of 15.6 over BLIP-2. It indicates that the local visual features and spatial information provided by the visual refiner effectively facilitate fine-grained visual-language alignment, thus improving the model’s ability to capture and respond to instruction-oriented visual objects. Furthermore, Table 2 presents our experiment results on text-oriented VQA benchmarks, from which we can observe that Lyrics significantly outperforms the latest Qwen-VL by 7.9 and 5.3 on the TextVQA and OCR-VQA benchmarks, and achieves comparable performance with PALI-X-55B [8], equipped with high-capacity image encoder and large language model. We believe that the improvement can be attributed to the introduction of semantic-aware visual objects extracted from MQ-Former, which facilitate the understanding of text within images.
5.3 Referring Expression Comprehension Results

### 12 Apr 2024 v2

@@ Abstract
Large Vision Language Models (LVLMs) have demonstrated impressive zero-shot capabilities in various vision-language dialogue scenarios. However, the absence of fine-grained visual object detection hinders the model from understanding the details of images, leading to irreparable visual hallucinations and factual errors. In this paper, we propose Lyrics, a novel multi-modal pre-training and instruction fine-tuning paradigm that bootstraps vision-language alignment from fine-grained cross-modal collaboration. Building on the foundation of BLIP-2, Lyrics infuses local visual features extracted from a visual refiner that includes image tagging, object detection and semantic segmentation modules into the Querying Transformer, while on the text side, the language inputs equip the boundary boxes and tags derived from the visual refiner. We further introduce a two-stage training scheme, in which the pre-training stage bridges the modality gap through explicit and comprehensive vision-language alignment targets. During the instruction fine-tuning stage, we introduce semantic-aware visual feature extraction, a crucial method that enables the model to extract informative features from concrete visual objects. Our approach achieves robust performance on 13 datasets across various vision-language tasks, and demonstrates promising multi-modal understanding, perception and conversation capabilities in 11 scenario-based benchmark toolkits.
@@ 1 Introduction
However, widely-used LVLMs habitually adopt Vision Transformer (ViT) [14] from pre-trained CLIP [47] as the image encoder, whose visual feature generalization capabilities are manifested in executing pre-defined label classification and brief image-text matching. Therefore, learning to effectively detect fine-grained visual objects within images (e.g. color, count, detailed description) and capturing visual morphology (e.g. action recognition, localization) present considerable challenges due to the lack of precise local visual features. Therefore, in situations where the image encoder fails to provide sufficient visual signals to meet the requirements of the specific objectives mentioned in the instructions, the LVLMs tends to produce incorrect responses that deviate from the details of the image. As analysed in Section 5.5., in dialogues involving visual objects, such as “How many people in the image?” and “What sport are the people playing?”, existing LVLMs are limited by visual signals and prone to generating visual hallucinations.
@@
we conduct extensive experiments on diverse vision-language tasks, including image captioning, visual question answering (VQA) and referring expression comprehension (REC). The results demonstrate that Lyrics can achieve state-of-the-art or comparable performance on several benchmarks compared to previous LVLMs.
@@ 2.2 Large Vision-Language Models
With remarkable generalization and robustness of LLMs, common LVLMs use a vision-language cross-modal adapter to align the visual features from the visual encoder with the LLMs, thereby stimulating the ability of LLMs to perceive and understand visual signals. Flamingo [2] freezes the pre-trained visual encoder and LLMs and integrates multi-modal representations through perceiver and gated cross-attention, demonstrating impressive few-shot capabilities. Meanwhile, BLIP-2 [29] trains a Q-Former to compress visual features as input to the frozen LLMs. On this basis, InstructBLIP [11] proposes instruction-aware visual feature extraction that enables flexible and informative feature extraction according to the given instructions. Early work such as LLaVA [31] and Mini-GPT4 [69] attempt to simply feed visual features into LLMs using only a learnable linear layer, which introduce visual instruction tuning to enhance instruction following capabilities in LVLMs. Furthermore, concurrent works such as Vision-LLM [58], Kosmos-2 [44], Shikra [6] and Qwen-VL [3] also demonstrate that the open training on visual encoders and LLMs can promote the LVLMs to understand located objects within the images and generate text formats of bounding boxes to perform visual grounding.
@@ 3.2 Bootstraping Vision-Language Representation Alignment via Multi-Task Pre-training
(ITM) is a binary classification task that transmits the pooled query embedding and into a classifier followed by softmax to predict a two-class probability and . It aims to learn image-text representations to capture both coarse-grained and fine-grained semantic alignments between vision and language. We employ a regionalized bi-directional self-attention mask that permits mutual interaction between visual queries and image caption, as well as between grounding queries and spatial representation, while remaining tokens are prevented from attaching to each other. Let and denote 2-dimension one-hot vectors representing the ground-truth label. The ITM loss is:
@@
The full training objective of MQ-Former can be formulated as:
@@ 4 Experiment Setting
Model | Image Captioning | General VQA | Text-oriented VQA | ||||||||||||||||||
VQAv2 | OKVQA | GQA |
Flamingo-9B | 79.4 | - | 61.5 | 51.8 | 44.7 | - | - | 28.8 | 31.8 | - | |||||||||||
Flamingo-80B | 84.3 | - | 67.2 | 56.3 | 50.6 | - | - | 31.6 | 35.0 | - | |||||||||||
IDEFICS-9B (LLaMA-7B) | 46.0 | 36.8 | 27.3 | 50.9 | 38.4 | - | 44.2 | 35.5 | 25.9 | - | |||||||||||
IDEFICS-80B (LLaMA-65B) | 91.8 | 65.0 | 53.7 | 60.0 | 45.2 | - | 68.9 | 36.0 | 30.9 | - | |||||||||||
BLIP-2 (Vicuna-13B) | - | 103.9 | 71.6 | 65.0 | 45.9 | 41.0 | 61.0 | 19.6 | 42.4 | - | |||||||||||
InstructBLIP (Vicuna-13B) | - | 121.9 | 82.8 | - | - | 49.5 | 63.1 | 33.4 | 50.7 | - | |||||||||||
Shikra (Vicuna-13B) | 117.5 | - | 73.9 | 77.4 | 47.2 | - | - | - | - | - | |||||||||||
Qwen-VL (Qwen-7B) | - | 120.2 | 81.0 | 78.2 | 56.6 | 57.5 | 68.2 | 38.9 | 61.5 | 70.5 | |||||||||||
Lyrics (Vicuna-13B) | 121.1 | 126.8 | 85.4 | 81.2 | 58.2 | 62.4 | 71.1 | 37.6 | 69.4 | 75.8 |
@@
Training Data
In the pre-training stage, we use a large-scale, web-crawled set of image-text pairs and filter out low-relevant samples. In the instruction fine-tuning stage, we first use a wide range of publicly available vision-language datasets and transform them into instruction fine-tuning format for multi-task learning. Then, we introduce high-quality vision-language annotation and instruction-response data to enhance the instruction following and dialogue capabilities of Lyrics. We present the detailed description and statistics for each dataset in Table 1.
Implementation Detail
For model settings, we choose ViT-L/14
[14] initialized from pre-trained CLIP [47] via contrastive learning as the image encoder. We build the visual refiner by combining Grounding-DINO-T [65] with 900 output object boxes and Swin-T backbone, SAM-HQ [25] with MAE and pre-trained VIT-H image encoder, and RAM++ [66] with Swin-B backbone. We use Vicuna-13B [9], an instruction-tuned variant from LLaMA [54], as the foundation backbone. Throughout the entire training process, the image encoder and visual refiner remains frozen. We focus on training the MQ-Former and linear projection layer, and efficient fine-tuning the large language model using LoRA [21]. With LoRA, we fine-tune the and via low-rank adaptation. We use images of size 224×224, augmented with random resized cropping and horizontal flipping.
We use AdamW [33] optimizer with , and a weight decay of 0.05. We use a cosine learning rate scheduler to train our model decay with a peak learning rate of 1e-4 and a linear warmup ratio of 15%. We train the Lyrics on 16xA100 GPUs for 800k steps in the vision-language representation alignment stage with a global batch size of 512, and 300k steps in the vision-to-language generative learning stage with a global batch size of 64.
@@ 5 Experiment Result
In this section, we conduct a comprehensive evaluation across various multi-modal tasks to thoroughly assess the visual understanding and generating capabilities of Lyrics, and compare our methods with the state-of-the-art visual-centric generalist models under zero-shot and few-shot settings, primarily including Flamingo [2], IDEFICS [27], LLaVA [31], BLIP-2 [29], InstructBLIP [11], Shikra [6], Qwen-VL [3], ShareGPT4V [7] and task-specific methods.
@@ 5.1 Dataset and Evaluation Metrics
We evaluate our model across a range of image captioning, VQA and REC benchmarks. For image captioning, we choose COCO [8], Nocaps [1] and Flickr30K [45] as benchmarks and report CIDEr score [56] as metric. We consider five benchmarks including VQAv2
[17], OKVQA [37], GQA [22], ScienceQA (Image Set) [34] and Vizwiz [19] benchmarks for general VQA, two benchmarks including TextVQA [51] and OCR-VQA [40] for text-oriented VQA, and evaluate the performance by matching the model’s response to the ground-truth and reporting top-1 accuracy. We use a sort of REC benchmarks such as RefCOCO [24], RefCOCO+ [36] and RefCOCOg [36] to verify the image understanding and localization capabilities. A predicted bounding box is considered as correct for reporting accuracy if its IOU between prediction and ground-truth is higher than 0.5. We use an open-ended approach with a greedy decoding strategy. We further conduct a comprehensive evaluation across 11 benchmark toolkits to thoroughly assess the multi-modal perception and conversation capabilities, which involve open-ended answers and factual assessments. Here we report the results in MathVista [35], MMMU [64], MME Perception (MME) [15], MME Cognition (MME) [15], MMBench (MMB) [32], MMBench-Chinese (MMB) [32], SEED-Bench Image Part
(SEED) [28], LLaVA-Bench In-the-Wild (LLaVA) [31], MM-Vet [63], QBench-Testset (QBench) [60] and HallusionBench (HallB) [18].
5.2 Image Understanding Results
Image Captioning and Visual Question Answering.
We first evaluate Lyrics on multiple image captioning and general VQA benchmarks. As demonstrated in Table 2, We discover that Lyrics achieves the best performance across 7 out of 8 benchmarks, and demonstrate competitive results in the remaining VizWiz. Lyrics consistently surpasses its original backbone BLIP-2 by a significant margin across all benchmarks, and achieves competitive performance to Qwen-VL, which possesses a more robust LLM backbone and underwent more pre-training and instruction fine-tuning steps. For instance, we achieve the 121.1, 126.8 and 85.4 state-of-the-art CIDEr scores on three image captioning benchmarks, even outperforms previous generalist models with much more parameters (e.g., Flamingo and IDEFICS with 80B parameters). We achieve 62.1 average accuracy on all benchmarks for general VQA tasks, representing a relative improvement of 15.6 over BLIP-2. It indicates that the local visual features and spatial information provided by the visual refiner effectively facilitate fine-grained visual-language alignment, thus improving the model’s ability to capture and respond to instruction-oriented visual objects. Furthermore, Table 2 also presents our experiment results on text-oriented VQA benchmarks, from which we can observe that Lyrics significantly outperforms the latest Qwen-VL by 7.9 and 5.3 on the TextVQA and OCR-VQA benchmarks. We believe that the improvement can be attributed to the introduction of semantic-aware visual objects extracted from MQ-Former, which facilitate the understanding of text within images.
Referring Expression Comprehension.
@@ 5.2 Image Understanding Results
Method | MathVista | MMMU | MME | MME | MMB | MMB | SEED | LLaVA | QBench | MM-Vet | HallB |
BLIP-2 (FLAN-T5) | - | 35.7 | 1293.8 | 290.0 | - | - | 46.4 | 38.1 | - | 22.4 | - |
InstructBLIP (Vicuna-7B) | 25.3 | 30.6 | - | - | 36.0 | 23.7 | 53.4 | 60.9 | 55.9 | 26.2 | 53.6 |
IDEFICS-80B (LLaMA-65B) | 26.2 | 24.0 | - | - | 54.5 | 38.1 | 52.0 | 56.9 | - | 39.7 | 46.1 |
Qwen-VL-Chat (Qwen-7B) | 33.8 | 35.9 | 1487.5 | 360.7 | 60.6 | 56.7 | 58.2 | 67.7 | 61.7 | 47.3 | 56.4 |
LLaVA (Vicuna-7B) | 23.7 | 32.3 | 807.0 | 247.9 | 34.1 | 14.1 | 25.5 | 63.0 | 54.7 | 26.7 | 44.1 |
LLaVA-1.5 (Vicuna-13B) | 26.1 | 36.4 | 1531.3 | 295.4 | 67.7 | 63.6 | 68.2 | 70.7 | 61.4 | 35.4 | 46.7 |
ShareGPT4V (Vicuna-7B) | 25.8 | 36.6 | 1567.4 | 376.4 | 68.8 | 62.2 | 69.7 | 72.6 | - | 37.6 | 49.8 |
Lyrics (Vicuna-13B) | 39.4 | 40.2 | 1597.3 | 431.6 | 75.3 | 62.4 | 71.8 | 76.9 | 72.5 | 46.3 | 62.6 |
5.3 Multi-Modal Benchmark Toolkit Results
In Table 4, we present a quantitative comparison between our proposed Lyrics model with existing SOTA LVLMs. Specifically, Lyrics outperforms the previously best-performing ShareGPT4V model by 3.6, 6.5 and 4.3 points on the MMMU, MMB, LLaVA benchmarks, demonstrating superior capabilities in tasks such as detailed description and complex reasoning. On the MME Benchmark, Lyrics achieves the highest scores in both perception (P) and cognition (C) capabilities, surpassing Qwen-VL-Chat in by 89.8 and 70.9 points, which was trained on 1.4 billion data. In the low-level image assessment QBench and multi-level image assessment with 14K questions SEED benchmarks, Lyrics achieves the highest score of 72.5% and 71.8%, 10.8% and 2.1% higher than the second-ranked LVLMs. which can be attributed to the diversity of our constructed dataset. Notably, Lyrics achieves a significant improvements on the MathVista and HallB benchmarks, demonstrating that the visual objects provided by the visual refiner can enhance the model’s capability to perceive real symbols and eliminate visual hallucinations.
Architecture | VQAv2 | RefCOCOg | LLaVA | HallB |
Lyrics | 81.2 | 88.26 | 76.9 | 62.6 |
w/o ViT | 78.5 | 86.11 | 74.6 | 60.1 |
w/o ODM | 80.6 | 87.55 | 76.1 | 60.7 |
w/o SSM | 80.8 | 87.08 | 76.5 | 59.2 |
w/o VR | 77.2 | 83.25 | 73.3 | 56.8 |